In three units of study, there will be application-focused cases due at the beginning of the class that will be provided by the instructor. These cases will be complex in nature and will require the application of course concepts to real-word business situations. Each case will have an associated rubric to highlight expectations. All submissions must be of professional quality and done in Microsoft Word, Microsoft Excel, or submitted as a PDF.


Objectives

To apply cost-volume profit analysis to a business decision
To calculate operating income and perform sensitivity analysis on operating income given changes in variables.
To illustrate the importance of hidden (undirected) issues that arise from a detailed analysis.
To prepare a coherent report and integrated analysis that meets specific user needs.

Instructions

In order to complete your case analysis successfully, you must

identify the role you are playing,
assess user needs
analyze user needs or issues (qualitatively and quantitatively), and
provide a recommendation and conclusion.

An average grade will result from answering all questions with basic coverage and accuracy, showing all your work. Additional points come from including greater detail, astute, informed commentary where appropriate and connections to readings and other content.

Case: Cost Structures for Global Shippers Inc.

Management from Global Shippers Inc, an international shipping business, is in the process of assessing the choice between two different cost structures for the business. Option A has relatively higher variable costs per unit shipped but lower annual fixed costs, while Option B has the opposite—relatively lower variable costs in its cost structure but higher fixed costs. Assume that delivery selling prices per unit are constant. The table below contains critical information in making the decision:

Cost Information

Option A

Option B

Delivery price (revenue) per shipment

$100

$100

Variable cost per shipment delivered

$85

$60

Contribution Margin per unit

$15

$40

Fixed costs (annual)

$1,450,000

$4,700,000


Management wants you to write a professional report, answering the following questions:

Questions

1) What is the break-even point, in terms of volume (i.e., number of shipments per year), for Option A? Option B?

(2) How many shipments would have to be made under Option A to produce operating income of $43,000 for an annual period?

(3) How many shipments per month would have to be made under Option A to produce an annual operating margin equal to 11% of sales revenue?

(4) How many shipments are required under Option B to produce net income of $213,000 per year, given a corporate tax rate of 25%?

(5) Assume that for the coming year total fixed costs are expected to decrease by 8% for each of the two options. What is the new break-even point, in terms of number of shipments, for each option? By what percentage did the break-even point change for each case? How do these figures compare to the percentage increase in budgeted fixed costs?

(6) Assume an average income-tax rate of 35%. What volume (number of shipments) would be needed to generate net income of 8% of revenue for each option?

(7) Which option do you think is the more profitable one for this business? Explain.

(8) Which option do you consider to be more risky to the business? Explain (calculate degree of operating leverage to help answer this question).
Answer #1

Solution

Global Shippers Inc

  1. Break-even point in terms of volume for

Option A –

Break-even point, in terms of volume = fixed cost/unit contribution margin

Fixed cost = $1,450,000

Contribution margin per unit = $15

Break-even points in terms of volume = 1,450,000/15 = 96,667 shipments

Option B –

Break-even point, in terms of volume = fixed cost/unit contribution margin

Fixed cost = $4,700,000

Contribution margin per unit = $40

Break-even points in terms of volume = 4,700,000/40 = 117,500 shipments

  1. Number of shipments needed –

Option A – to produce operating income of $43,000 for an annual period:

Desired number of shipments = (target income + fixed cost)/unit contribution margin

Target income = $43,000

Fixed cost = $1,450,000

Unit contribution margin = $15

Desired number of shipments = (43,000 + 1,450,000)/15 = 99,533 shipments

  1. Number of shipments needed to earn operating margin equal to 11% of sales revenue:

Option A – to produce operating margin equal to 11% of sales revenue.

Desired number of shipments = (target income + fixed cost)/unit contribution margin

Assuming the desired number of shipments to be A,

Target income = 11% x $100 x A = 11A

Fixed cost = $1,450,000

Unit contribution margin = $15

Desired number of shipments = (11A + 1,450,000)/15 = A shipments

15A = 11A + 1,450,000

4A = 1,450,000

A = 1,450,000/4 = 362,500 shipments

Target income = 11% x $100 x 362,500 = $3,987,500

  1. Number of shipments required under Option B to produce net income of $213,000 per year, given a corporate tax rate of 25%:

Net income =213,000

Tax rate = 25%

Before tax income = 213,000/75% = $284,000

Fixed cost = $4,700,000

Contribution margin per unit = $40

Number of shipments = (284,000 +4,700,000)/40 = 124,600 shipments

Hence, shipments needed to produce after tax net income of $213,000 per year under option B = 124,600.

  1. Determination of new break-even point when fixed cost decrease by 8% for both the options:

Option A –

Fixed cost = $1,450,000

Less: Decrease = 8% x 1,450,000 = 116,000

New fixed cost = 1,334,000

Contribution margin per unit = $15

Break-even point, in terms of number of shipments = 1,334,000/15 = 88,933 shipments

Change in break-even point –

Original break-even point 96,667 shipments

New break-even point in terms of volume = 88,933

decrease = 96,667 – 88,933 = 7,734 shipments

Percent decrease in break-even point, volume = 7,734/96,667 = 8%

Option B –

Fixed cost = $4,700,000

Less: decrease = 8% x 4,700,000 = 376,000

New fixed cost = 4,324,000

Contribution margin per unit = $40

Break-even point, in terms of number of shipments = 4,324,000/40 = 108,100 shipments

Original beak-even point in volume = 117,500 shipments

New break-even point , volume = 108,100 shipments

decrease = 117,500 – 108,100 = 9,400 shipments

Percent decrease = 9,400/117,500 = 8%

The decrease in break-even point, in terms of shipment for both the options is 8%, which is equal to the percent decrease in fixed cost for both the options. This indicates that the break-even point, in terms of value increases in proportionately with the fixed cost.

  1. Shipments needed to generate net income of 8% revenue for each option, assuming a tax rate of 35%:

Option A -

Net income = 8% x $100 x A = $8A

Tax rate = 35%

Before tax income = 8A/65% = $12.31A

Fixed cost = $1,450,000

Contribution margin per unit = $15

Number of shipments = (12.31A +1,450,000)/15 = A

15A = 12.31A +1,450,000

2.69 A = 1,450,000

A = 1,450,000/2.69 = 539,033 shipments

Option B –

Net income = 8% x $100 x A = $8A

Tax rate = 35%

Before tax income = 8A/65% = $12.31A

Fixed cost = $4,700,000

Contribution margin per unit = $40

Number of shipments = (12.31A + 4,700,000)/40 = A

40A = 12.31A +4,700,000

27.69 A = 4,700,000

A = 4,700,000/27.69 = 169,736 shipments

Hence, shipments needed to produce net income of 8% of sales revenue per year under option B = 169,736.

  1. The most profitable options:

Option A is the most profitable option for this business. The break-even point, in terms of volume for option A (96,667) is less compared to the break-even point, in terms of volume for option B (117,500). The lesser the break-even point, the quicker the company would start making profits.

  1. More risky option –

The option B is probably the risky option, in view of higher proportion of fixed costs.